Suppose each firm's long run average cost curve, for positive levels of output, is given by AC = 0.1 + 0.05Q + 5/Q. The marginal cost curve is given by MC = 0.1 + 0.1Q.

(a) Find the minimum efficient scale for the above cost function.

(b) What is the firm's minimum average cost?

(c) Suppose you have many identical firms in a long run competitive equilibrium. Demand is P = 13.1-0.04Q. What is the market quantity? How many firms are there?

(d) Suppose demand increases to P = 15.1- 0.04Q. Supposing that increases in industry output do not affect input prices, what is the new long run competitive

equilibrium, and how many new firms will there be? Illustrate your new and old competitive equilibrium on a graph with the demand curves and the long run supply curve?
Answer #1

AC= 1t.058 +5/9, a) Min AC So dAC 5-5 0 92 .05=5/92 92-X 100 So =10Min efecent Scale fn b) Now Min AC = 1t -05xIO+ 5110 1+ 5